Status Report

Hubble Space Telescope Daily Report 2774 12/21/00

By SpaceRef Editor
December 21, 2000
Filed under ,

PERIOD COVERED:  0000Z (UTC) 12/20/00 – 0000Z (UTC) 12/21/00
Daily Status Report as of 356/0000Z
    1.1 Completed Two Sets of WF/PC-2 8601 (A Snapshot Survey of Probable Nearby Galaxies)
        The WF/PC-2 was used to continue the very successful snapshot survey in order to use the high spatial resolution of HST to determine whether selected galaxies are nearby on the basis of resolution into stars, and the magnitudes and colors of the brightest stars.  There were no reported problems.
    1.2 Completed Five Sets of STIS/CCD 8260 (Searching for the Hydrogen Reionization Edge of the Universe at 5HST)
        The Space Telescope Imaging Spectrograph (CCD) was used to observe one of 36 parallel orbits {4–5 fields of 5–8 orbits each} to constrain the H Lyman-edge in emission that marks the transition from a neutral to a fully ionized IGM at a predicted zion~eq5–15.  This edge is due to recombination from the H Lyman series and Lyman continuum, and can be used to constrain zion, one of the most important unknown quantities in large scale structure and cosmology.  The proposal completed nominally
    1.3 Completed Two Sets of WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3)
        The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels.  The proposal completed nominally.
    1.4 Completed Four Sets of STIS/CCD 8836 (Performance Monitor {Cycle 9})
        The Space Telescope Imaging Spectrograph (CCD) was used to measure the baseline performance and commandability of the CCD subsystem.  Only primary amplifier D was used.  Bias and Flat Field
exposures were taken in order to measure read noise, CTE, and
gain.  Numerous bias frames were taken to permit construction of "superbias" frames in which the effects of read noise have been rendered negligible.  Full frame and binned observations were made, with binning factors of 2 x 1, 1 x 2, 2 x 2, 4 x 1, and 4 x 2.  Dark images were taken in 2×2 binning mode; 1×1 binning darks are being taken in the nominal CCD Dark Monitor.  The exercise completed nominally.
    1.5 Completed Two Sets of STIS/CCD 8837 (CCD Dark Monitor-Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks.  The proposal completed with no reported problems.
    1.6 Completed STIS/CCD/MA2 8111 (CS 22892–052: A Rosetta Star for the Age and Early History of the Galaxy)
        The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of the "Rosetta" star CS 22892-052.  These observations should provide significant new information about the early history of the galaxy, specifically the nature of the earliest generations of stars and the types of nucleosynthetic processes that occurred in those stars.  The observations were executed as scheduled, and no anomalies were noted.
    1.7 Completed Five Sets of WF/PC-2 8883 (POMS Test Proposal: WFII backup parallel archive proposal II)
        The WF/PC-2 was used to complete a POMS test proposal designed to simulate scientific plans.  There were no reported problems.
    1.8 Completed Four Sets of WF/PC-2 8815 (Cycle 9 Earth Flats)
        The WF/PC-2 was used to monitor flatfield stability.  This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set.  These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats.  The proposal completed nominally.
    1.9 Completed STIS/CCD 8838 (Bias Monitor – Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.  The proposal completed nominally.
    1.10 Completed FGS/1R 8834 (Calibrating Interferometric Response as a Function of Spectral Color)
        Fine Guidance Sensor 1R was used to obtain reference point source Transfer Functions {S-Curves} through the F583W filter at the center position of the FOV for a variety of spectral colors.  The proposal completed with no reported problems.
    1.11 Completed STIS/MA2 8843 (Cycle 9 MAMA Dark Measurements)
        The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise.  The proposal completed nominally.
    1.12 Completed FGS/1R 8469 (Calibrating FGS1R’s Optical Field Angle Distortion {OFAD})
        Fine Guidance Sensor 1R was used to gather the data needed to calibrate the optical field angle distortions in FGS1R to the level of accuracy required for astometry science.  Selected stars from the galactic cluster M35 are repeatedly observed in POSITION mode by FGS1R with the F583W filter at a variety of spacecraft pointings and telescope roll angles.  There were no problems.
    1.13 Completed Two Sets of STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations)
        The Space Telescope Imaging Spectrograph (CCD) was used to probe the distortion of light bundles from distant galaxies, looking at the statistical properties of the intervening inhomogeneous {dark} matter distribution.  The proposal completed nominally.
    1.14 Completed WF/PC-2 8584 (Calibrating the Metallicity Dependence of the Cepheid PL Relation)
        The WF/PC-2 was used to determine the uncertainty in the metal abundance dependence of the Cepheid PL relation that remains as one of the largest sources of systematic error in the Cepheid distance scale and in the extragalactic distance scale as a whole.  There were no reported problems.
    1.15 Completed WF/PC-2 8669 (Merger-Driven Evolution Of Galactic Nuclei: Observations Of The Toomre Sequence)
        The WF/PC-2 was used to observe galaxy mergers that are believed responsible for triggering starburst and AGN activity in galaxies, and even perhaps transforming spiral galaxies into ellipticals.  The proposal completed nominally.
    2.1 Guide Star Acquisitions:
      Scheduled Acquisitions: 10
Successful: 10
    Scheduled Re-acquisitions:  5
Successful:  5
    2.2 FHST Updates:
                  Scheduled: 17
Successful: 14
        Per HSTAR 8014, the full maneuver updates scheduled at 356/075535Z and at 356/075820Z failed due to FHST #1.  The subsequent guide star acquisition, however, was successful.
        As described in HSTAR 8015, the full maneuver update scheduled for 356/094148Z failed due to tracker #1.  The vehicle attitude was very close to what it had been for the previous failures.
    2.3 Operations Notes:
        Using ROP SR-1A, the SSR EDAC error counter was cleared three times.
        The engineering status buffer limits were adjusted three times per ROP DF-18A.
        The NSSC-1 status buffer was dumped and reset at 355/1527Z as directed by ROP NS-3.
        Continuation of normal science observations and calibrations.

SpaceRef staff editor.